OpenStudy (anonymous):

Vector B (with arrow above) has x, y, and z components of 6.00, 8.00, and 5.00 units, respectively. Calculate the magnitude of B (with arrow above). ______ Calulcate the angle that makes with the x axis. ____° Calculate the angle that makes with the y axis. ____° Calculate the angle that makes with the z axis. ____°

OpenStudy (anonymous):

\[\left| B \right|=\sqrt{6^{2}+8^{2}+5^{2}}\]

OpenStudy (kira_yamato):

\[|B| = \sqrt (6^2+8^2+5^2) = 5 \sqrt 5\] \[\cos \theta _{x}= 6/\sqrt125\] \[\cos \theta _{y}= 8/\sqrt125\] \[\cos \theta _{z}= 5/\sqrt125\] Please calculate the angles yourself by taking the arcosine of each angle on these equations as I prefer to leave things in the exact form. Just punch your calculate to find that stuff...